OpenStudy (lovelyanna):

Use the discriminant to determine the number and type of solutions the equation has. x^2 + 6x + 12 = 0 A. one real solution B. two rational solutions C. no real solution D. two irrational solutions

OpenStudy (lovelyanna):

@jigglypuff314

OpenStudy (lovelyanna):

I don't think it's A or B

OpenStudy (anonymous):

x2 + 2x – 2 = 0 Discriminant = b^2 - 4ac, a = 1, b = 2, c = - 2 = (2)^2 - 4(1)(- 2) = 4 + 8 = 12 D)

OpenStudy (lovelyanna):

So it's D? Thought so! Thanks :)

jigglypuff314 (jigglypuff314):

I thought it was C... http://www.wolframalpha.com/input/?i=x%5E2+%2B+6x+%2B+12+%3D+0+

OpenStudy (lovelyanna):

I can't read that o.O Can you explain in your own words, Emily? XP

jigglypuff314 (jigglypuff314):

it does not touch the x axis so there are no real solutions because solutions come from when the graph touches the x axis

OpenStudy (lovelyanna):

Mmm, confusing... o.O

jigglypuff314 (jigglypuff314):

oh, jordanloveangle wrote a -2 instead of a 12

OpenStudy (lovelyanna):

Sooooo...?

jigglypuff314 (jigglypuff314):

x2 + 2x + 12 = 0 Discriminant = b^2 - 4ac, a = 1, b = 2, c = 12 = (2)^2 - 4(1)(12) = 4 - 48 = negative -> means no real solutions http://www.mathwarehouse.com/quadratic/discriminant-in-quadratic-equation.php

OpenStudy (lovelyanna):

Mmm, makes sense :) I agree with you

OpenStudy (lovelyanna):

Emily, how about this one? What number would you have to add to both sides to complete the square? x^2 – 14x = 7 A. –49 B. 196 C. –196 D. 49 I think D.

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (lovelyanna):

Yesh :D Thanks :)